If a negative charge is placed in an electric field, will t move toward a region of higher potential or lower potential? What will happen to its electric potential energy? What will happen to its kinetic energy? What will happen to its velocity? For a particles with charge Q, what is the formula for the Electric Potential, V, and the formula for the Electric Field, E? Electrons in a TV tube are accelerated from rest by a potential difference of 1000 V. If the potential difference is increased to 2000 V then how much greater will their velocity be?
Answer #1

a)

electron moves from higher to lower potential.

electric potential energy decreases.

using concept of energy conservation, decreasing lectric potential energy will increase kinetic energy of electron.

increase in kinetic energy will increase the velocity of electron.

=====

b)

let us consider d is the distnace from the charge where we need to find electric potential or field

Electric potential, V = k Q / d

electric field E = k Q / d^2

=======

c)

as we know

0.5 m v^2 = qV

so, square of velocity is directly proportional to potential difference

v1^2 / v2^2 = V1 / V2

v2 / v1 = sqrt ( V2 / V1)

v2 / v1 = sqrt (2000/1000)

v2 / v1 = sqrt (2)

so new velocity must be sqrt(2) timea of previous value.

======

do comment in case any doubt, will reply for sure.. Goodluck